# Methodology, Investigation, Software, Visualization, Writing – original draft

### v1

@@ 1 Introduction
Therefore, we hypothesize that conventional timestamp automatic annotation, which rely heavily on high-quality timestamp labels to yield satisfactory performance, may experience decreased performance when applied to datasets with noisy timestamp labels, and current approach aimed at addressing this issue rely on probability sampling, which may result in the selection of unimportant samples. To mitigate these challenges, we developed a novel timestamp calibration method for time-series ScRNA-seq data, referred to as ScPace. In contrast to previous method that correct cell type labels using semi-supervised Adasampling techniques, our approach introduces a latent variable indicator for the selection of potentially mislabeled samples, enabling their deletion or reclassification according to specific requirements. This method is informed by the principles of self-paced learning[12], a machine learning training paradigm that emulates the cognitive processes involved in human learning by progressively transitioning from easier to more difficult samples, which has being applied to various fields like multi-modal image classification[13], domain adaption[14], imbalance sampling[15] and functional brain network scrubbing[16]. Specially, previous research on imbalance sampling utilize self paced learning has achieved better performance than random imbalance sampling methods, which emphasizes the retention of high-quality samples with accurate label information while discarding those identified as potential mislabeled points.
@@
The following passages are organized as follows. In Section 2, we introduced some related work. In Section 3, our proposed method ScPace is described here. In Section 4 the experimental design and experiment results of cross validation with different mislabeling rate and enhancement of supervised pseudotime analysis are shown, moreover, we also assessed the detected mislabeled cells in simulated and real time-series ScRNA-seq data. The final discussion are in Section 5 and conclusion in Section6.
@@ 3.1.1 Preprocessing of Dimensional Reduction
3.1.2 Optimizing One-Vs-One binary classifier with latent variable
In this section, we will introduce the methodology for training a single weighted One-vs-One binary Support Vector Machine (SVM) classifier with the introduced latent variable. This latent variable will be further optimized in subsequent sections. The One-vs-One approach allows us to construct multiple binary classifiers for each pair of classes, enabling effective discrimination even in multi-class scenarios. Since SVM adopts One-vs-One strategies to deal with multi-class problem, therefore $y=1$ are assigned to one classes and $y=-1$ are assigned to the distinct class. Given the reduced or original training sets $T=\{(x_{1},y_{1}),(x_{2},y_{2}),\dots,(x_{n},y_{n})\}$ where $x_{i}\in\{R^{m^{{}^{\prime}}},R^{m}\},i=1,2,\dots,n$, $y\in\{-1,+1\}$, The standard soft-margin SVM classifier is being optimized with the primal problem equation:
@@ 3.1.2 Optimizing One-Vs-One binary classifier with latent variable
Moreover, The rewritten equation could be further transformed into hinge loss format which is equivalent to optimizing the primal problem, therefore Eq.(4) would then rewrite into Eq.(6):
@@
Since the dual problem is a maximizing optimization problem on the minima of Lagrangian functions. Therefore, we must seek partial derivatives of $w$,$b$,$\zeta_{i}$ of Eq.(7). Then, we get the following equations:
@@ 3.1.3 Optimize latent variable with optimized parameters $w$ and $b$
For multi-class problem, SVM initially adopts the ove-vs-one training strategy, which means that $\frac{N(N-1)}{2}$ binary classifiers are being trained in Section(3.1.2), and $N$ represents number of classes. After Computing the parameter $w$ and $b$ according to the dual coefficients, we therefore could obtain $\frac{N(N-1)}{2}$ Decision Values for each samples according to Eq.(15). Here we transformed the decision value shape to One-Vs-Rest form using a transformed confidence strategy where the decision value of each samples would be $N$ dimension instead of $\frac{N(N-1)}{2}$ dimension. The algorithm are written according to Scikit-learn[34].
@@
According to Algorithm 1, we therefore could obtain the voting matrix $\hat{V}$ and the sum confidence matrix $S$, we then monotonically transform $S$ according to the following Equation:
$\hat{S}=\frac{S}{3\left|S\right|+3}$ (17)
The final decision value transformed into One-vs-Rest Shape are therefore the sum of voting matrix and transformed sum confidence matrix $\hat{S}$.
@@ 4.1 Data Preparation
We used the R package Splatter[36] to simulate two types of trajectories: a linear path and a bifurcation path[37]. For the linear path, designated as Sim1, we simulated 1,500 cells, each characterized by 6,000 genes. Each cell state corresponds to a single time point, resulting in four different time points, each with ordinal labels ($T_{1},T_{2},T_{3},T_{4}$). Next, we simulated two bifurcation paths, one balanced and one imbalanced. The balanced data, referred to as Sim2, consists of 1,500 cells and 8,000 genes, with three time points ($T_{1},T_{2},T_{3}$). The imbalanced data, referred to as Sim3, includes 2,000 cells and 10,000 genes, featuring four time points ($T_{1},T_{2},T_{3},T_{4}$). For the balanced bifurcation path in Sim2, the cell state prior to the split into two branches is defined as a single time point. Here, State 1 (S1) is considered as $T_{1}$. In Sim3, the same labeling is applied, with S1 as $T_{1}$ and State 2 (S2) as $T_{2}$. After the split in Sim2, both State 3 and State 4 are assigned to $T_{2}$, while States 5 and 6 are assigned to $T_{3}$. In Sim3, States 3 and 4 are considered as $T_{3}$, and States 5 and 6 correspond to $T_{4}$. We also simulated dropouts, which is a common feature in ScRNA-seq datasets. Subsequently, we conducted gene filtering on the simulated datasets to enhance the data quality for further analysis.
For real time-series ScRNA-seq datasets, all datasets were publicly downloaded from the US National Library of Medicine, National Institutes of Health. For GSE90047 and GSE98664[38], we first log-normalized them using the Python Scanpy[39] package and removed genes that were expressed in fewer than 10 cells. Cells with low counts were also excluded if the total number of genes expressed in a cell was lower than 200. However, given that the original dataset for GSE98664 comprised 157,717 genes, we subsequently selected only the highly variable genes for further analysis using Scanpy. In the case of GSE67310[40], the dataset was already normalized upon download, therefore, our preprocessing focused on gene and cell filtering only. Additional statistical information regarding both the simulated and real time-series ScRNA-seq datasets is presented in Table 1.
@@
Table 1: Statistical results on Simulated Datasets and real time-series ScRNA-seq datasets, each rows correspond to different datasets, Sim1 and Sim2 are class balanced datasets with different trajectory, and Sim3 correspond to imbalanced dataset on bifurcation path. Genes and cells are listed in terms of the preprocessed datasets.
@@ 4.2.1 Experimental Setup
We conducted stratified k-fold(k=5) cross validation on the simulated datasets and real time-series ScRNA-seq datasets. k-fold stratified cross validation is a variant of the k-fold cross-validation method that aims to ensure that the class distribution in the dataset is maintained across all folds. The term "stratified" means that the division is done in a way that preserves the original class distribution of the data. This is particularly useful when dealing with imbalanced datasets, where the number of samples in different classes varies significantly. In our cross-validation experiments, the training sets are artificially mislabeled and the holdout testing sets are cleaned without artificial mislabeling. When assigning the artificial mislabeling on the training sets, there are two types of mislabeling.
  • The first type of mislabeling we consider is known as swap mislabeling. In this method, cells are randomly selected from the training sets based on a specified mislabeling rate (e.g., 0.1, 0.2, 0.3, or 0.4). Each selected cell is assigned a decision value of either 00 or $1$: $1$ indicates that the cell will be mislabeled, while 00 signifies that the cell remains clean (i.e., correctly labeled). For the cells marked for mislabeling, their original labels are swapped with labels from neighboring time points. For cells labeled at intermediate time points $T_{m}$ where $T_{1}<T_{m}<T_{n}$, the selected cells are mislabeled as either $T_{m-1}$ or $T_{m+1}$. Specially, cells labeled as $T_{1}$ would be mislabeled as $T_{2}$, for cells labeled with $T_{n}$, they would instead be mislabeled as $T_{n-1}$. This structured approach to mislabeling helps to maintain some degree of temporal coherence in the dataset, as the labels are swapped only with those from neighboring time points. This mimics potential real-world scenarios where labels may become inaccurate but remain somewhat related to nearby observations in time-series ScRNA-seq datasets.
@@
we compared our proposed method against several traditional machine learning classifiers that do not specifically account for noise in the data. For this comparison, we included two tree-based classifiers: Decision Tree and Random Forest. Additionally, we incorporated a deep learning approach using a fully connected neural network classifier. It is important to note that ScReclassify is not originally designed for predicting labels for new cells. To address this limitation, we extended the methodology by performing dimensional reduction on the testing set, using the parameters determined from the training set.
We employed the grid search method for tuning the hyperparameters of the models to comprehensively compare the effectiveness of each classifier. Particularly, special attention was given while conducting the cross-validation experiments on the fully connected neural network, as it is prone to extreme overfitting. The parameters set for the grid search are presented in Table 2:
@@
[…]
{(50,100,50),(100,100,50)}
@@ 4.2.2 Experimental Results
In comparison to ScReclassify, ScPace consistently demonstrated superior performance across all mislabeling rates and both types of mislabeling. Notably, as the mislabeling rate increased, ScPace maintained strong performance levels, whereas the performance of ScReclassify exhibited a decline. This highlights ScPace’s resilience and reliability, particularly in challenging conditions where mislabeling is prevalent. Overall, the results affirm ScPace’s effectiveness as a robust classifier across various situations.
@@
On real time-series ScRNA-seq datasets, we also introduced two types of mislabeling for cross-validation experiments. The experimental results shown in Fig.(3) that on ScPace performed better than ScReclassify and traditional machine learning classifiers on most occassions, even when mislabeling rate reaches $40\%$. Interestingly, we found out that when mislabeling rate is under 40%, the performance of SVM tends to perform better than ScReclassify. Interestingly, we found out that when there is no artificial mislabeling, all the classifier performed above accuracy score $95\%$, which indicates that the label noises in the testing sets didn’t effect the overall performance, practically because the timestamps of most cells in the original are well labeled.
@@
4.2.3 Sensitivity Analysis
4.2.4 Assessing the performance of Dimensional Reduction
@@ 4.3.1 Experimental Setup
In addition to cross-validation, another reliable way to assess the effectiveness of our proposed method is to evaluate whether the calibrated timestamps can enhance the performance of supervised pseudotime analysis. Previous research has shown that supervised pseudotime analysis is sensitive to the quality of timestamps. To investigate this further, we compared the enhancement results of our method against those achieved using ScReclassify for calibrating timestamps across both simulated and original time-series ScRNA-seq datasets. Note that ScPace includes two types of timestamp calibration: deletion and reclassification. In contrast, ScReclassify focuses exclusively on reclassification. Therefore, when comparing our method with ScReclassify, we restricted our analysis specifically to the reclassification of noisy simulated and real time-series ScRNA-seq datasets.
We began our analysis by performing reclassification using both ScPace and ScReclassify on noisy simulated datasets with two types of artificial mislabeling, and then applied the reclassification on original real time-series ScRNA-seq datasets without artificial mislabeling rate.
After completing the reclassification using both ScPace and ScReclassify on the original noisy time-series ScRNA-seq data, we applied Psupertime to the reclassified timestamps from both methods independently and then computed the average training accuracy, testing accuracy and pseudotime value to each cells. To assess the performance of the computed pseudotime, we calculated the Spearman and Kendall correlation between the computed pseudotime value and the calibrated timestamps. Then we applied psupertime to the original timestamps for comparison purpose, seeking whether the calibration of timestamps could enhance the performance.
We also implemented the deletion calibration method of ScPace and compared the results with original timestamps. Through these comparisons, we aimed to evaluate the overall performance of our proposed method in enhancing the results of supervised pseudotime analysis.

### v2

@@ 1 Introduction
Therefore, we hypothesize that conventional timestamp automatic annotation, which rely heavily on high-quality timestamp labels to yield satisfactory performance, may experience decreased performance when applied to datasets with noisy timestamp labels, and current approach aimed at addressing this issue rely on probability sampling, which may result in the selection of unimportant samples. To mitigate these challenges, we developed a novel timestamp calibration method for time-series ScRNA-seq data, referred to as ScPace. In contrast to previous method that correct cell type labels using semi-supervised Adasampling techniques, our approach introduces a latent variable indicator for the selection of potentially mislabeled samples, enabling their deletion or reclassification according to specific requirements. This method is informed by the principles of self-paced learning[12], a machine learning training paradigm that emulates the cognitive processes involved in human learning by progressively transitioning from easier to more difficult samples, which has being applied to various fields like multi-modal image classification[13], domain adaption[14], imbalance sampling[15] and functional brain network scrubbing[16]. Specially, previous research on imbalance sampling utilize self paced learning has achieved better performance than random imbalance sampling methods, which emphasizes the retention of high-quality samples while discarding bad-quality samples to construct a balanced samples.
@@
The following passages are organized as follows. In Section 2, we introduced some related work. In Section 3, our proposed method ScPace is described here. In Section 4 the experimental design and experiment results of cross validation with different mislabeling rate and enhancement of supervised pseudotime analysis are shown. Moreover, We also included two case studies to illustrate the practical implications of using ScPace in real-world applications in Section 5, and the computational timing of ScPace in Section 6. The final discussion are in Section 7 and conclusion in Section 8.
@@ 3.1.1 Preprocessing of Dimensional Reduction
3.1.2 Optimize One-Vs-One binary classifiers with latent variable
In this section, we will introduce the methodology for training a single weighted One-vs-One binary Support Vector Machine (SVM) classifier with the introduced latent variable. This latent variable will be further optimized in subsequent sections. The One-vs-One approach allows us to construct multiple binary classifiers for each pair of classes, enabling effective discrimination even in multi-class scenarios. In binary cases, therefore $y=1$ are assigned to one classes and $y=-1$ are assigned to the distinct class. Given the reduced or original training sets $T=\{(x_{1},y_{1}),(x_{2},y_{2}),\dots,(x_{n},y_{n})\}$ where $x_{i}\in\{R^{m^{{}^{\prime}}},R^{m}\},i=1,2,\dots,n$, $y\in\{-1,+1\}$, The standard soft-margin SVM classifier is being optimized with the primal problem equation:
@@ 3.1.2 Optimize One-Vs-One binary classifiers with latent variable
Moreover, The rewritten equation could be further transformed into hinge loss format which is equivalent to optimizing the primal problem shown in Eq.(6).
@@
Since the dual problem is a maximizing optimization problem on the minima of Lagrangian functions. Therefore, we must seek partial derivatives of $w$,$b$,$\zeta_{i}$ of Eq.(7).
@@ 3.1.3 Optimize latent variable with optimized parameters $w$ and $b$
For multi-class problem, SVM initially adopts the ove-vs-one training strategy. After Computing the parameter $w$ and $b$ in terms of the dual coefficients, we therefore could obtain $\frac{N(N-1)}{2}$ Decision Values for each samples according to Eq.(15). Here we transformed the decision value shape into $N$ dimensions using a transformed confidence strategy.
@@
In terms of Algorithm 1, we therefore could obtain the voting matrix $\hat{V}$ and the sum confidence matrix $S$, we then monotonically transform $S$ according to the following Equation:
$\hat{a_{ij}}=\frac{a_{ij}}{3(\left|a_{ij}\right|+1)}$ (17)
Where $\hat{a_{ij}}\in\hat{S}$, $a_{ij}\in S$, ($i=1,2,\dots,n$, $j=1,2,\dots,N$). The final decision value transformed into One-vs-Rest Shape are therefore the sum of voting matrix and transformed sum confidence matrix $\hat{S}$.
@@ 4.1 Data Preparation
We used the R package Splatter[35] to simulate two types of trajectories: a linear path and a bifurcation path[36]. For the linear path, designated as Sim1, we simulated 1,200 cells, each characterized by 6,000 genes. Each cell state corresponds to a single time point, resulting in four different time points, each with ordinal labels ($T_{1},T_{2},T_{3},T_{4}$). Next, we simulated two bifurcation paths, one balanced and one imbalanced. The balanced data, referred to as Sim2, consists of 1,500 cells and 8,000 genes, with three time points ($T_{1},T_{2},T_{3}$). The imbalanced data, referred to as Sim3, includes 2,000 cells and 10,000 genes, featuring four time points ($T_{1},T_{2},T_{3},T_{4}$). For the balanced bifurcation path in Sim2, the cell state prior to the split into two branches is defined as a single time point. Here, State 1 (S1) is considered as $T_{1}$. In Sim3, the same labeling is applied, with S1 as $T_{1}$ and State 2 (S2) as $T_{2}$. After the split in Sim2, both State 3 and State 4 are assigned to $T_{2}$, while States 5 and 6 are assigned to $T_{3}$. In Sim3, States 3 and 4 are considered as $T_{3}$, and States 5 and 6 correspond to $T_{4}$. We also simulated dropouts, which is a common feature in ScRNA-seq datasets. Subsequently, we conducted gene filtering on the simulated datasets to enhance the data quality for further analysis.The visualization of the simulated datasets are shown in Fig.(2).
For real time-series ScRNA-seq datasets, all datasets were publicly downloaded from the US National Library of Medicine, National Institutes of Health. For GSE90047 and GSE98664[37], we first log-normalized them using the Python Scanpy[38] package and removed genes that were expressed in fewer than 10 cells. Cells with low counts were also excluded if the total number of genes expressed in a cell was lower than 200. However, given that the original dataset for GSE98664 comprised 157,717 genes, we subsequently selected only the highly variable genes for further analysis using Scanpy. In the case of GSE67310[39], the dataset was already normalized upon download, therefore, our preprocessing focused on gene and cell filtering only. For dataset E-MTAB-3929, low quality cells and genes are first removed and then we only selected highly variable genes for further analysis. Additional statistical information regarding both the simulated and real time-series ScRNA-seq datasets is presented in Table 1.
@@
      E-MTAB-3929       1529       10235       5       F
Table 1: Statistical results on Simulated Datasets and real time-series ScRNA-seq datasets, each rows correspond to different datasets, Sim1 and Sim2 are class balanced datasets with different trajectory, linear and bifurcation, and Sim3 correspond to imbalanced dataset on bifurcation path. Genes and cells are listed in terms of the preprocessed datasets.
@@ 4.2.1 Experimental Setup
We conducted stratified k-fold(k=5) cross validation on the simulated datasets and real time-series ScRNA-seq datasets. k-fold stratified cross validation is a variant of the k-fold cross-validation method that aims to ensure that the class distribution in the dataset is maintained across all folds. The term "stratified" means that the division is done in a way that preserves the original class distribution of the data. This is particularly useful when dealing with imbalanced datasets, where the number of samples in different classes varies significantly. In our cross-validation experiments, the training sets are artificially mislabeled and the holdout testing sets are cleaned without artificial mislabeling. When assigning the artificial mislabeling on the training sets, there are two types of mislabeling which are shown in Fig.(3).
  • The first type of mislabeling we consider is known as swap mislabeling. In this method, cells are randomly selected from the training sets based on a specified mislabeling rate (e.g., 0.1, 0.2, 0.3, or 0.4). Each cell is assigned a decision value of either 00 or $1$. In this case, $1$ indicates that the cell will be mislabeled, while 00 signifies that the cell remains clean (i.e., correctly labeled). For the cells marked for mislabeling, their original labels are swapped with labels from neighboring time points. For cells labeled at intermediate time points $T_{m}$ where $T_{1}<T_{m}<T_{n}$, the selected cells are mislabeled as either $T_{m-1}$ or $T_{m+1}$. Specially, cells labeled as $T_{1}$ would be mislabeled as $T_{2}$, for cells labeled with $T_{n}$, they would instead be mislabeled as $T_{n-1}$. This structured approach to mislabeling helps to maintain some degree of temporal coherence in the dataset, as the labels are swapped only with those from neighboring time points. This mimics potential real-world scenarios where labels may become inaccurate but remain somewhat related to nearby observations in time-series ScRNA-seq datasets.
@@
we compared our proposed method against several traditional machine learning classifiers that do not specifically account for noise in the data. For this comparison, we included two tree-based classifiers: Decision Tree and Random Forest. Additionally, we incorporated a deep learning approach using a fully connected neural network classifier. It is important to note that ScReclassify is not originally designed for predicting labels for new cells. To address this limitation, we extended the methodology by performing dimensional reduction on the testing set, using the parameters determined from the training set. Moreover, we have included fully connected neural network based confident learning[40] using CleanLab for timestamp calibration.
For the setting of the model hyperparameters, we utilized the grid search method to systematically tune the hyperparameters of the models, thereby enabling a comprehensive comparison of the effectiveness of each classifier. The grid search method is an exhaustive approach for hyperparameter tuning in machine learning models, involving a thorough search through a manually specified subset of the hyperparameter space to identify the optimal combination that yields the best model performance on a validation set. Particularly, special attention was given while conducting the cross-validation experiments on the fully connected neural network, as it is prone to extreme overfitting. The parameters set for the grid search are presented in Table 2:
@@
Neural Network (i,j,k) The hidden layer sizes of the network {(50,50,50),(50,100,50)}
@@
CL (i,j,k) The hidden layer sizes of the network {(50,50,50),(50,50,100)}
$\rho(x)$ The activation function of each neurons $\max{(0,x)}$}
@@
Moreover, we also assessed the correction ability of ScPace, ScReclassify and CL using Eq.(25). This metric assess the amount of mislabeled samples corrected compared with the number of all mislabeling samples. Given mislabeling rate $\rho$, the artificial mislabeled samples $\{(x_{1},y_{1}),\dots,(x_{\rho n},y_{\rho n})\}$, the cleaned samples $\{(x_{1},y_{1}^{*}),\dots,(x_{\rho n},y_{\rho n}^{*})\}$ and the reclassified samples $\{(x_{1},\hat{y_{1}}),\dots,(x_{\rho n},\hat{y_{\rho n}})\}$:
$C=\frac{\sum_{i=1}^{\rho n}1(y^{*}_{i}=\hat{y_{i}})}{\rho n}$ (25)
@@ 4.2.2 Experimental Results
On real time-series ScRNA-seq datasets, we also introduced two types of mislabeling for cross-validation experiments. The experimental results shown in Fig.(5) that on ScPace performed better than ScReclassify and traditional machine learning classifiers on most occasions, even when mislabeling rate reaches $40\%$. Interestingly, we found out that when mislabeling rate is under 40%, the performance of SVM tends to perform better than ScReclassify. Interestingly, we found out that when there is no artificial mislabeling, all the classifier performed above accuracy score $95\%$, which indicates that the label noises in the testing sets didn’t effect the overall performance, practically because the timestamps of most cells in the original are well labeled.
In comparison to ScReclassify and CL, ScPace consistently demonstrated superior performance across all mislabeling rates and both types of mislabeling. Notably, as the mislabeling rate increased, ScPace maintained strong performance levels, whereas the performance of ScReclassify and CL exhibited a decline. The correction score of the three calibration methods shown in Fig.(6) for simulated datasets and Fig.(7) for real time-series ScRNA-seq demonstrate the superior correction performance of ScPace on most occasions. All the results highlights ScPace’s resilience and reliability, particularly in challenging conditions where mislabeling is prevalent. Overall, the results affirm ScPace’s effectiveness as a robust classifier across various situations.
@@
4.2.3 Sensitive Analysis
Improper tuning of hyperparameters can lead to erroneous results in ScPace. To evaluate the impact of hyperparameter variation on model performance, we conducted a sensitivity analysis by systematically adjusting key parameters. These included $C$, the regularization term; $p$, the C-Growing parameter; iter, the number of iterations for running ScPace; $\lambda$, the threshold for updating latent variables. During the analysis of each parameter, the selected hyperparameter was varied while holding others constant. Additionally, the dimensionality reduction method was fixed to PCA throughout the sensitivity analysis.
The results Shown in Fig.(8) for simulated datasets and Fig.(9) for real time-series ScRNA-seq datasets that besides hyperparameter $C$, the performance of those hyperparameter, including $\lambda$,$Iter$,and $p$, doesn’t show significant varying performance. In contrast, we observed that there is a significant varying performance on hyperparameter C on most dataset.
Interestingly, We observed that if the parameter $C$ is set too low or too high, ScPace tends to perform adequately. This sensitivity to hyperparameter tuning highlights the need for a balanced approach to parameter selection, ensuring that the calibration process optimally identifies and mitigates mislabeled cells without unnecessarily compromising the dataset’s integrity.
@@ 4.3.1 Experimental Setup
In addition to cross-validation, another reliable way to assess the effectiveness of our proposed method is to evaluate whether the calibrated timestamps can enhance the performance of supervised pseudotime analysis. Previous research has shown that supervised pseudotime analysis is sensitive to the quality of timestamps. To investigate this further, we compared the enhancement results of our method against those achieved using ScReclassify for calibrating timestamps across both simulated and original time-series ScRNA-seq datasets. Note that both ScPace and CL includes two types of timestamp calibration: deletion and reclassification. In contrast, ScReclassify focuses exclusively on reclassification. Therefore, when comparing our method with ScReclassify, we restricted our analysis specifically to the reclassification of noisy simulated and real time-series ScRNA-seq datasets.
We initiated our analysis by applying reclassification through three calibration methods to both the artificially mislabeled simulated datasets and the original real-time series ScRNA-seq datasets (which were not artificially mislabeled).
After completing the reclassification using the three calibration methods on the original noisy simulated and real time-series ScRNA-seq data, we applied Psupertime to the reclassified timestamps from both methods independently and then computed the average training accuracy, testing accuracy and pseudotime value to each cells. To assess the performance of the computed pseudotime, we calculated the Spearman and Kendall correlation between the computed pseudotime value and the calibrated timestamps. Then we applied psupertime to the original timestamps for comparison purpose, seeking whether the calibration of timestamps could enhance the performance.
We also implemented the deletion calibration method of ScPace and CL and compared the results with original timestamps. Through these comparisons, we aimed to evaluate the overall performance of our proposed method in enhancing the results of supervised pseudotime analysis.
@@
CL (i,j,k) The hidden layer sizes of the network {(50,50,50),(50,50,100)}
$\rho(x)$ The activation function of each neurons $\max{(0,x)}$}